# Efficient Testing Implies Structured Symmetry

### v1

@@ Abstract
Given a small random sample of $n$-bit strings labeled by an unknown Boolean function, which properties of this function can be tested computationally efficiently? We show an equivalence between properties that are efficiently testable from few samples and properties with structured symmetry, which depend only on the function’s average values on parts of a low-complexity partition of the domain. Without the efficiency constraint, a similar characterization in terms of unstructured symmetry was obtained by Blais and Yoshida (2019). Our main technical tool is supersimulation, which builds on methods from the algorithmic fairness literature to approximate arbitrarily complex functions by small-circuit simulators that fool significantly larger distinguishers.
We extend the characterization along other axes as well. We show that allowing parts to overlap exponentially reduces their required number, broadening the scope of the construction from properties testable with $O(\log n)$ samples to properties testable with $O(n)$ samples. For larger sample sizes, we show that any efficient tester is essentially checking for indistinguishability from a bounded collection of small circuits, in the spirit of a characterization of testable graph properties. Finally, we show that our results for Boolean function testing generalize to high-entropy distribution testing on arbitrary domains.
@@ 1 Introduction
The goal in distribution testing [GGR96] is to determine whether an unknown probability distribution $\mathcal{D}$ has a particular property of interest, such as uniformity, independence, or equality to a reference distribution. Rather than being given a complete description of $\mathcal{D}$, the algorithm has access to an oracle that generates independent samples from $\mathcal{D}$ on request. Ideally, we would like to make as few requests as possible, run a fairly simple computation on the resulting samples, and determine whether $\mathcal{D}$ has the property or is far from having the property.
In this work, we focus on distributions over pairs $(x,y)\in\{0,1\}^{n}\times\{0,1\}$, where $x$ is a uniform $n$-bit string, and $y=f(x)$ for some unknown Boolean function $f:\{0,1\}^{n}\to\{0,1\}$. Our main result is a characterization of which properties of $f$ can be efficiently tested from few samples in this framework. By “efficient,” we mean that the testing algorithm can be represented as a small Boolean circuit that takes as input a set of labeled samples and outputs Accept or Reject.
@@ 1.1 Background
Without the requirement of computational efficiency, an elegant result of Blais and Yoshida [BY19] fully characterizes the properties of Boolean functions that are testable from a constant number of samples (i.e. not scaling with $n$). They showed that a property is constant-sample testable if and only if it has constant-part symmetry, where a property of $f$ is said to have $k$-part symmetry if it depends only on the average values of $f$ within some fixed partition of the domain into parts $S_{1},\ldots,S_{k}\subseteq\{0,1\}^{n}$. Equivalently, permutations of the domain within each part do not affect whether or not $f$ has the property.
@@
Theorem 1.1 gives a transformation from $\mathcal{P}$ to $\mathcal{Q}$ that loses computational efficiency (ours will not). Indeed, $\mathcal{Q}$ may be computationally intractable even when $\mathcal{P}$ has a computationally efficient tester. To illustrate this, consider any $k$-part symmetric property $\mathcal{Q}$ with invariant sets $S_{1},\ldots,S_{k}$. Clearly, $\mathcal{Q}$ can be tested from a handful of $(x_{i},y_{i})$ samples—just classify each $x_{i}$ according to the part $S_{j}$ that contains it, and use the corresponding labels $y_{i}$ to empirically estimate the average value of $f$ over $S_{j}$. Given these $k$ averages, testing for $\mathcal{Q}$ can always be done efficiently, simply because any function on few inputs can be computed with a small circuit by brute force. However, the overall procedure may still be computationally inefficient—indeed, because the sets $S_{j}$ are unstructured, performing even a single classification could require up to $\exp(n)$ time!
@@ 1.2 Results Overview
Consider the case that $m$, $k$, and $\varepsilon$ are constants, so that the circuit size $s$ is the only parameter varying with $n$. In this case, Theorem 1.1 states that constant-sample testability is equivalent to constant-part symmetry, while Theorems 1.2 and 1.3 state that constant-sample testability with a circuit of size $O(s)$ is equivalent to constant-part symmetry with partition complexity $O(s)$.
To prove Theorems 1.2 and 1.3, our main technical tool is supersimulation [DT25], which builds on a series of constructive generalizations of the complexity-theoretic regularity lemma [TTV09] pioneered by [HKRR18] in the algorithmic fairness literature (see also [DKR+21, GKR+22, CDV24], which are more closely related). The basic versions of these results take labeled samples of the form $(x_{i},f(x_{i}))$ and a function class $\mathcal{F}$ as input, and build a simulator $\tilde{f}$ making few oracle calls to members of $\mathcal{F}$ and fooling distinguishers in $\mathcal{F}$. Supersimulators, however, also fool distinguishers that are significantly more powerful than themselves.
Our proof first simulates the oracle that generates the samples, and then simulates the tester itself. Interestingly, between the two steps, the roles of the simulators and distinguishers switch. We discuss our proof strategy further in the technical overview in Section 1.3, and give a full proof in Section 3.
Extensions
In Section 4, we establish various other results regarding property testing of Boolean functions from samples, beyond our main results (Theorems 1.2 and 1.3).
First, in Section 4.1, we address another drawback of the result of [BY19]: the doubly exponential dependence on the tester’s sample complexity $m$ in the number of parts of the partition $S_{1},\ldots,S_{k}$. We show that allowing parts to overlap can exponentially improve this dependence, reducing the number of structured sets from $2^{2^{O(m)}}$ to just $2^{O(m)}$. We actually show an even stronger result: If $\mathcal{P}$ is $O(m)$-sample testable, then there exists a tester for $\mathcal{P}$ which simply counts the number of functions $f\in\mathcal{F}_{\mathrm{good}}$ and $f\in\mathcal{F}_{\mathrm{bad}}$ that are consistent with its $O(m)$ observed samples—the tester outputs Accept if these samples are consistent with more good functions than bad, and Reject otherwise. Here, $\mathcal{F}_{\mathrm{good}}$ and $\mathcal{F}_{\mathrm{bad}}$ are some prespecified families of at most $2^{O(m)}$ Boolean functions, which depend on the property $\mathcal{P}$. For this result, we do not consider their computational complexity. From another point of view, a $k$-part partition of a domain of size $2^{n}$ is nontrivial only if $k<2^{n}$. Thus, constructing an $2^{2^{O(m)}}$-part partition is only interesting for properties testable with $m=O(\log n)$ samples, but our theorem remains meaningful for a broader class of properties, namely those testable with up to $O(n)$ samples.
@@
Next, in Section 4.2, we consider properties that are testable using $m$ samples and a circuit of size $s$, where we only require that $m$ and $s$ are subexponential in $n$. In this setting, we turn our attention from $k$-part symmetry to regularity templates, in the spirit of a landmark characterization of testable graph properties, due to [AFNS06]. In that work, it was shown that a property of dense graphs is testable from a constant number of edge queries if and only if having the property is roughly equivalent to having a regular partition from some finite set of templates. Here, a “regular partition” refers to the kind provided by the Szemerédi regularity lemma [Sze75]: a partition of the graph’s vertices into a small number of parts between which the graph’s edges are distributed pseudorandomly. By analogy, our result shows that if a property of Boolean functions is testable using $m$ samples and a circuit of size $s$, then having the property is roughly equivalent to being indistinguishable from some circuit of size $O(m^{2}s)$ in a bounded collection, where indistinguishability is defined with respect to circuits of size $s$.
Finally, in Section 4.3, we sketch a straightforward argument that our main results extend beyond the setting of testing Boolean functions from samples. Specifically, we can apply our results to distribution testing over arbitrary domains, provided that those distributions have sufficiently high min-entropy.

### v2

# Efficient and Private Property Testing via Indistinguishability

@@ Abstract
Given a small random sample of $n$-bit strings labeled by an unknown Boolean function, which properties of this function can be tested computationally efficiently? We show an equivalence between properties that are efficiently testable from few samples and properties with structured symmetry, which depend only on the function’s average values on an efficiently computable partition of the domain. Without the efficiency constraint, a similar characterization in terms of unstructured symmetry was obtained by Blais and Yoshida (2019). We also give a function testing analogue of the classic characterization of testable graph properties in terms of regular partitions, as well as a sublinear time and differentially private algorithm to compute concise summaries of such partitions of graphs. Finally, we tighten a recent characterization of the computational indistinguishability of product distributions, which encompasses the related task of efficiently testing which of two candidate functions labeled the observed samples.
Essential to our proofs is the following observation of independent interest: Every randomized Boolean function, no matter how complex, admits a supersimulator: a randomized polynomial-size circuit whose output on random inputs cannot be efficiently distinguished from reality with constant advantage, even by polynomially larger distinguishers. This surprising fact is implicit in a theorem of Dwork et al. (2021) in the context of algorithmic fairness, but its complexity-theoretic implications were not previously explored. We give a new proof of this lemma using an iteration technique from the graph regularity literature, and we observe that a subtle quantifier switch allows it to powerfully circumvent known barriers to improving the landmark complexity-theoretic regularity lemma of Trevisan, Tulsiani, and Vadhan (2009).
@@ 1 Introduction
The goal in distribution testing [GGR96, BFR+00] is to determine whether an unknown probability distribution $\mathcal{D}$ has a particular property of interest, such as uniformity or equality to a reference distribution. Rather than being given a complete description of $\mathcal{D}$, the algorithm has access to an oracle that generates independent samples from $\mathcal{D}$ on request. Ideally, we would like to make as few requests as possible, run a fairly simple computation on the resulting samples, and determine whether $\mathcal{D}$ has the property or is far from having the property.
In this work, we focus on distributions over pairs $(x,y)\in\{0,1\}^{n}\times\{0,1\}$, where $x$ is a uniform $n$-bit string, and $y=f(x)$ for some unknown Boolean function $f:\{0,1\}^{n}\to\{0,1\}$. Our main result is a characterization of which properties of $f$ can be efficiently tested from few samples in this framework. By “efficient,” we mean that the testing algorithm can be represented as a small Boolean circuit that takes as input a set of labeled samples and outputs Accept or Reject. We also design a sublinear time differentially private [DMNS06] algorithm to compute concise summaries of regular partitions of graphs, which are known to characterize testability for graph properties [AFNS06], and give a function testing version of the characterization. Finally, we tighten a recent characterization of the computational indistinguishability of product distributions, which corresponds to the conceptually easier task of simple hypothesis testing.
@@ 1.1 Background
Without the requirement of computational efficiency, an elegant result of Blais and Yoshida fully characterizes the properties of Boolean functions that are testable from a constant number of samples (i.e. not scaling with $n$) [BY19]. They showed that a property is constant-sample testable if and only if it has constant-part symmetry, where a property of $f$ is said to have $k$-part symmetry if permutations of the domain within parts of some fixed partition $S_{1},\ldots,S_{k}\subseteq\{0,1\}^{n}$ do not affect whether or not $f$ has the property. Equivalently, the property depends only on the average values of $f$ within each part.
@@
Theorem 1.1 gives a transformation from $\mathcal{P}$ to $\mathcal{Q}$ that loses computational efficiency (ours will not). Indeed, $\mathcal{Q}$ may be computationally intractable to test even when $\mathcal{P}$ has a computationally efficient tester. To illustrate this, consider any $k$-part symmetric property $\mathcal{Q}$ with label-invariant sets $S_{1},\ldots,S_{k}$. Clearly, $\mathcal{Q}$ can be tested from a handful of $(x_{i},y_{i})$ samples—just classify each $x_{i}$ according to the part $S_{j}$ that contains it, and use the corresponding labels $y_{i}$ to empirically estimate the sum of $f$ over $S_{j}$. Given these $k$ sums, testing for $\mathcal{Q}$ can always be done efficiently, simply because any function on few inputs can be computed with a small circuit by brute force. However, the overall procedure may still be computationally inefficient—indeed, because the sets $S_{j}$ are unstructured, performing even a single classification could require up to $\exp(n)$ time!
@@ 1.2 Results Overview
We first strengthen the forward direction of Theorem 1.1. While Theorem 1.1 does not bound the partition complexity of $\mathcal{Q}$, which may be as large as $\exp(n)$, the following theorem bounds the partition complexity by $2^{O(m)}s$, which is a potentially much smaller quantity.
@@
When combined, Theorems 1.2 and 1.3 give a characterization of properties efficiently testable from a constant number of samples. To illustrate this, suppose that $m$, $k$, and $\varepsilon$ are constants, so that the circuit size $s$ is the only parameter varying with $n$. In this case, Theorem 1.1 states that testability is equivalent to symmetry, while our Theorems 1.2 and 1.3 state that testability with a circuit of size $O(s)$ is equivalent to symmetry with partition complexity $O(s)$. We prove these theorems in Section 3. Of the two, Theorem 1.2 is the harder to prove. Its proof uses a tool that we call supersimulators, which we discuss further in the technical overview in Section 1.3. At a high level, the tool is a strengthening of the complexity-theoretic regularity lemma of [TTV09] that arose implicitly in the algorithmic fairness literature. Thus, Theorem 1.2 and its proof can be viewed as situated at the intersection of property testing, pseudorandomness (specifically, the structure-vs-randomness paradigm exemplified by the regularity lemma), and algorithmic fairness.
Theorems 1.1, 1.2, and 1.3 characterize not only constant-sample testability but also, for example, testability with $m(n)$ samples where $m(n)=O(\log^{(k)}(n))$ for all $k\in\mathbb{N}$ and $2^{n}$ is the domain size. Here, $\log^{(k)}$ denotes the $k$-fold iterated logarithm, so this class includes, for example, the function $m(n)=\log^{*}(n)$. In fact, even for properties testable from as many as $O(\log n)$ samples, the more interesting direction of our equivalence—efficient testing implies structured symmetry—continues to hold. Our next result pushes further in this direction, shedding light on the structure of properties that may require even more than $O(\log n)$ samples to test. Specifically, we present a modification to Theorem 1.1 that circumvents the $2^{2^{O(m)}}$ dependence on the tester’s sample size $m$, which is clearly a prerequisite for making any meaningful statement about sample sizes larger than $O(\log n)$. Specifically, we will show in Section 3.4 that any $m$-sample testable property essentially boils down to what we call consistency counting on a collection of $2^{O(m)}$ functions.
Definition 1.4 (Consistency Counting).
We say that a collection of $m$ samples $(x_{i},y_{i})$ is consistent with a function $f$ if the label $y_{i}=f(x_{i})$ for all indices $i\in[m]$. Given a family of good functions $\mathcal{F}_{+}$ and a family of bad functions $\mathcal{F}_{-}$, the $(m,\mathcal{F}_{+},\mathcal{F}_{-})$-consistency counter is the tester that receives $m$ inputs and outputs Accept iff they are consistent with strictly more good functions than bad.
Consistency counters are especially simple testers, which just check whether each of the prespecified functions in $\mathcal{F}_{+}\cup\mathcal{F}_{-}$ could have plausibly labeled the $m$ observed samples. The following theorem states that every property that is testable from few samples has a consistency counter using a similar number of samples. From another point of view, it gives a two-sided error analogue of the following folklore fact regarding one-sided sample-based property testing: If a property is testable with one-sided error, then it is testable with the “canonical” one-sided tester which exhaustively checks whether all observed samples are consistent with at least one function with the property.
Theorem 1.5.
There is a straightforward converse to Theorem 1.5: By definition, the $(m,\mathcal{F}_{+},\mathcal{F}_{-})$-consistency counter uses just $m$ samples, so if a property is testable with such a consistency counter, then it is testable with $m$ samples.
So far, we have framed our discussion around sample-testing Boolean functions, but in some sense, this framing is more restrictive than necessary. We remark that our main result, Theorem 1.2, can be straightforwardly extended to properties of distributions on an arbitrary domain $\mathcal{X}$, provided that the distribution to be tested is known in advance to have high density or, equivalently, high min-entropy. We state the extension in terms of $\mu$-densely testable properties, whose formal definition we defer to Section 3.5.
Theorem 1.6.
Let $0<\mu\leq 1/2$. If a distribution property $\mathcal{P}$ is $\mu$-densely testable with proximity $\varepsilon$ using $m$ samples and a circuit of size $s$, then $\mathcal{P}\subseteq\mathcal{Q}\subseteq\mathcal{P}_{\varepsilon}$ for some $2^{(1/\mu)^{O(m)}}$-part symmetric distribution property $\mathcal{Q}$ with partition complexity $(1/\mu)^{O(m)}s$.
1.2.1 Graph Property Testing
@@
Our next several results are motivated by the classic characterization of testable graph properties of Alon et al. [AFNS06]. In that work, it was shown that a graph property is testable from few edge queries if and only if the property is roughly equivalent to having a regular partition compatible with one of finitely many prespecified templates. A “regular partition” refers to the kind provided by the Szemerédi regularity lemma: a partition of the graph’s vertices into a small number of parts between which the graph’s edges are distributed pseudorandomly [Sze75]. Roughly speaking, a “template” is represented by a list of the pairwise densities between pairs of such parts.
Our first result in this setting is an extremely efficient (sublinear time) node-level differentially private algorithm to compute templates corresponding to such regular partitions. Although our algorithm is extremely natural, we will need to leverage nontrivial results from [AFNS06] for its analysis. Our second result in this setting will be an analogue of the characterization for efficiently testable Boolean function properties, as opposed to graph properties.
To state our results on private graph testing, we must first review some key notation regarding regular partitions of graphs. For convenience, we focus on simple graphs $G=(V,E)$. Given vertex sets $A,B\subseteq V$, we write $e_{G}(A,B)=\lvert E\cap(A\times B)\rvert$ for the number of edges from $A$ to $B$. We write $d_{G}(A,B)=e_{G}(A,B)/\lvert A\rvert\lvert B\rvert$ for the edge density from $A$ to $B$. We omit the subscript $G$ from $e_{G}$ and $d_{G}$ when clear from context. A partition $(V_{1},\ldots,V_{k})$ of the vertices $V$ is an equipartition if any two parts $V_{i},V_{j}$ differ in size by at most $1$. We denote the total number of vertices by $n=\lvert V\rvert$, and we write $G[Q]$ for the subgraph of $G$ induced by a vertex subset $Q\subseteq V$.
Definition 1.7 (Graph Regularity [Sze75]).
A pair of vertex sets $(A,B)$ in a graph $G=(V,E)$ is $\gamma$-regular if for all subsets $A^{\prime}\subseteq A$ and $B^{\prime}\subseteq B$ satisfying $\lvert A^{\prime}\rvert\geq\gamma\lvert A\rvert$ and $\lvert B^{\prime}\rvert\geq\gamma\lvert B\rvert$,
$\bigl\lvert d(A^{\prime},B^{\prime})-d(A,B)\bigr\rvert\leq\gamma.$
An equipartition $(V_{1},\ldots,V_{k})$ is $\gamma$-regular if at most $\gamma k^{2}$ of the pairs $(V_{i},V_{j})$ fail to be $\gamma$-regular.
Theorem 1.8 ([Sze75]).
Every graph has a $\gamma$-regular equipartition of size $k=O_{\gamma}(1)$. [^2] The suppressed constant is a power tower of $2$s of height $\mathrm{poly}(1/\gamma)$, which is large but does not scale with $n$.
The main result of [AFNS06] is a characterization in terms of regular partitions of testable properties in the dense graph model. In this model, properties are required to be invariant under vertex relabeling, and the tester accesses the graph via edge membership queries (i.e. “is an edge present between these vertices $u$ and vv?”). The distance between two graphs on the same vertex set is the fraction of edges on which they differ, or, more precisely, their edit distance under edge additions/deletions, divided by $n^{2}$. The characterization of testability is stated in terms of regularity templates and regular-reducibility:
Definition 1.9 (Regularity Template [AFNS06]).
A regularity template $R$ is defined by parameters $\gamma\in(0,1]$ and $k\in\mathbb{N}$, a function $\eta:[k]^{2}\to[0,1]$, and a set $\bar{R}\subseteq[k]^{2}$ of size at most $\gamma k^{2}$. A graph $G$ satisfies $R$ if it has an equipartition $(V_{1},\ldots,V_{k})$ such that $d(V_{i},V_{j})=\eta_{ij}$ for all pairs $(i,j)\in[k]^{2}$ and $(V_{i},V_{j})$ is $\gamma$-regular for all pairs $(i,j)\notin\bar{R}$. [^3] As in [AFNS06], when we say $d(V_{i},V_{j})=\eta_{ij}$, we really mean that $e(V_{i},V_{j})=\lfloor\eta_{ij}\lvert V_{i}\rvert\lvert V_{j}\rvert\rfloor$. We say that $G$ is $\varepsilon$-close to satisfying $R$ if $G$ is $\varepsilon$-close to some graph $G^{\prime}$ satisfying $R$. The complexity of $R$ is $\max(k,1/\gamma)$.
Definition 1.10 (Regular-Reducible [AFNS06]).
A property $\mathcal{P}$ of $n$-vertex graphs is regular-reducible if for all $\varepsilon>0$, there exists a set $\mathcal{R}$ of $O_{\varepsilon}(1)$ templates, each of complexity $O_{\varepsilon}(1)$, such that $\mathcal{P}\subseteq\mathcal{Q}_{\varepsilon}$ and $\mathcal{Q}\subseteq\mathcal{P}_{\varepsilon}$, where $\mathcal{Q}$ denotes the property of satisfying some template in $\mathcal{R}$. [^4] The choice of which regularity templates belong to the set $\mathcal{R}$ is allowed to depend on $n$, as long the number of templates and their complexity remain bounded by a constant $O_{\varepsilon}(1)$ independent of nn. [^5] The condition $\mathcal{P}\subseteq\mathcal{Q}_{\varepsilon}$ and $\mathcal{Q}\subseteq\mathcal{P}_{\varepsilon}$” is slightly weaker than the condition $\mathcal{P}\subseteq\mathcal{Q}\subseteq\mathcal{P}_{\varepsilon}$ that appears elsewhere in this paper. The latter ensures that any tester for $\mathcal{Q}$ is also a tester for $\mathcal{P}$. The former only ensures that any tolerant tester for $\mathcal{Q}$ is also a tester for $\mathcal{P}$. In the case of graph property testing, however, tolerant and non-tolerant testing are known to be roughly equivalent [FN05, GKS23].
Theorem 1.11 ([AFNS06]).
A property $\mathcal{P}$ of $n$-vertex graphs is regular-reducible if and only if for all $\varepsilon>0$, the property $\mathcal{P}$ is testable with proximity $\varepsilon$ using $O_{\varepsilon}(1)$ edge queries.
Given the importance of regular partitions of graphs, a natural question is whether such summaries can be computed efficiently, ideally in time sublinear in the number of vertices $n$. When data privacy is not a concern, the answer turns out to be yes: one particularly elegant and simple algorithm is to take a uniformly random sample of “anchor” vertices, fully reveal their neighborhoods, and use the intersections of these neighborhoods to partition the graph [Tao09]. However, this approach may be unsatisfactory in real-world networks, where privacy is a concern and the presence or absence of edges between nodes may be sensitive information. Our first result in this section is a modified algorithm for efficiently computing regularity templates while satisfying differential privacy (DP) [DMNS06]. In fact, we go beyond edge privacy to the challenging setting of node-level differential privacy, whose formal definition we defer to Section 2.
Theorem 1.12.
For all $\gamma,\varepsilon>0$, there exist parameters $k,q=O_{\gamma,\varepsilon}(1)$ and $\lambda=\Omega_{\gamma,\varepsilon}(1)$ such that for all input graphs $G$ on $n$ vertices, with probability at least $3/5$, Algorithm 1 with these parameters outputs a regularity template $R$ with $\gamma$ error and $k$ parts such that $G$ is $\varepsilon$-close to satisfying $R$. Moreover, Algorithm 1 satisfies node-level $\alpha$-differential privacy for $\alpha=O_{\gamma,\varepsilon}(1/n)$.
Algorithm 1 Node-Level Differentially Private Regular Partition
1:input: $G=(V,E)$ and $k,q\in\mathbb{N}$ and $\gamma,\lambda\in(0,1)$
[…]
4:for all $R\in\mathcal{R}$ do
[…]
7:return $R^{\prime}$
Although we defer the full analysis of Algorithm 1 to Section 4, the idea behind it is very simple. First, we take a random subsample of $q$ vertices from the graph, and we perform all $\binom{q}{2}$ edge queries between pairs of these vertices to obtain the induced subgraph $G[Q]$. Then, we exhaustively search over all possible regularity templates of a fixed granularity. We assign each candidate template $R$ a score depending on how close $G[Q]$ is to satisfying $R$, and then sample one such $R$ from a probability distribution weighted according to these scores. This is the classic exponential mechanism [MT07] from the DP literature.
Ultimately, since the number of templates is independent of $n$, our algorithm runs in constant $O_{\gamma,\varepsilon}(1)$ time, not counting the time required to sample $q=O_{\gamma,\varepsilon}(1)$ vertices from $V$. The fact that analyzing such a small number of vertices suffices to determine the partition structure of the overall graph is highly nontrivial, and it will require us to leverage results from [AFNS06]. We emphasize that what makes Algorithm 1 interesting is its efficiency. For example, if one were concerned only with privacy and approximation quality, but not efficiency, one could simply post-process any state-of-the-art private cut sparsifier, which are known to be computable in $\mathrm{poly}(n)$ time with error subsumed by that of the regularity construction [GRU12, AU19, EKKL20, LUZ24, ZLU25, ACD+25]. In contrast, our algorithm takes sublinear time.
We also extend the characterization of testability of Theorem 1.11 from the non-private to the private setting. The proof of the following theorem, which appears in Section 4, is based on a reduction from the private case to the non-private case, although it can also be proved directly using techniques similar to our analysis of Algorithm 1.
Theorem 1.13.
A property $\mathcal{P}$ of $n$-vertex graphs is regular-reducible if and only if for all $\varepsilon>0$, the property $\mathcal{P}$ is testable with proximity $\varepsilon$ using $O_{\varepsilon}(1)$ edge queries and node-level $O_{\varepsilon}(1/n)$-DP.
To conclude this section, we briefly shift our attention from graph testing back to function testing, with the goal of proving a function testing analogue of the (original, non-private) characterization from [AFNS06]. In this setting, we show that if a property of Boolean functions is testable using $m$ samples and a Boolean circuit of size $s$, then the property is essentially assessing compatibility with a collection $\mathcal{T}$ of appropriately defined regularity templates, each of complexity $O(m^{2}s)$. In particular, this result is nontrivial for all arbitrary sample sizes $m$ and circuit sizes $s$, as long as they are subexponential in $n$. In the Boolean circuit context, we shall model a template by a simulator for the function to be tested, as per the complexity-theoretic regularity lemma of [TTV09]. For us, compatibility will be the usual notion of computational indistinguishability. [^6] We say two functions $g,h$ are $(s,\delta)$-indistinguishable if $\lvert\operatorname*{\mathbb{E}}[f(x)(g(x)-h(x))]\rvert\leq\delta$ for all functions $f$ of complexity at most $s$. See Definition 2.2 for more details.
Definition 1.14.
Given a family $\mathcal{T}\subseteq\{\{0,1\}^{n}\to[0,1]\}$ (the templates), the compatibility property $\mathcal{T}_{s,\delta}$ is the set of Boolean functions that are $(s,\delta)$-indistinguishable from some function in $\mathcal{T}$.
The following result states that small-circuit testability, even for a moderate or large number of samples, reduces to testing compatibility with a size-bounded collection of regularity templates.
Theorem 1.15.
If $\mathcal{P}$ is testable with proximity $\varepsilon$ using $m$ samples and a circuit of size $s$, then there exists a set $\mathcal{T}$ of templates, each of circuit complexity $O(m^{2}s)$, such that $\mathcal{P}\subseteq\mathcal{T}_{s,\frac{1}{13m}}\subseteq\mathcal{P}_{\varepsilon}$.
We also prove a partial converse to the preceding theorem. We say “partial” because it only preserves sample efficiency, not computational efficiency.
Theorem 1.16.
Let $s\leq t$ with $t\log(t)\leq O(\varepsilon^{2}\delta^{2}2^{n})$. If $\mathcal{T}$ is a set of templates, each of circuit complexity at most $t$, then $\mathcal{T}_{s,\delta}$ is testable with proximity $\varepsilon$ from $O(t\log(t)/\varepsilon^{2}\delta^{2})$ samples.
1.2.2 Simple Hypothesis Testing
All of the function testing results discussed so far concern the broad task of distinguishing a property $\mathcal{P}$ from the complement of $\mathcal{P}_{\varepsilon}$, both of which may be doubly exponentially large sets of Boolean functions. Building on the recent work of [MPV25], our next few results concern the narrower task of simple hypothesis testing, which requires distinguishing between just two candidates.
For context, consider the following slightly more general question: Given $k$ independent samples from a distribution $\mathcal{D}_{b}\in\{\mathcal{D}_{0},\mathcal{D}_{1}\}$ over $\{0,1\}^{n}$, where $b\in\{0,1\}$ is an unknown bit, what is the best distinguishing advantage between $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$ that one could hope to efficiently achieve? With a computationally unbounded distinguisher, the answer is the total variation distance between the $k$-fold products of $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$, denoted $d_{\mathrm{TV}}(\mathcal{D}_{0}^{\otimes k},\mathcal{D}_{1}^{\otimes k})$. It is easy to show that this quantity relates to $d_{\mathrm{TV}}(\mathcal{D}_{0},\mathcal{D}_{1})$ via the following standard inequalities:
$d_{\mathrm{TV}}(\mathcal{D}_{0}^{\otimes k},\mathcal{D}_{1}^{\otimes k})\leq 1-(1-d_{\mathrm{TV}}(\mathcal{D}_{0},\mathcal{D}_{1}))^{k}\leq k\cdot d_{\mathrm{TV}}(\mathcal{D}_{0},\mathcal{D}_{1}).$
If we let $d_{s}(\cdot,\cdot)$ denote the best distinguishing advantage attainable by a circuit of size at most $s$, then it turns out that proving the analogous relationship between $d_{s}(\mathcal{D}_{0}^{\otimes k},\mathcal{D}_{1}^{\otimes k})$ and $d_{s}(\mathcal{D}_{0},\mathcal{D}_{1})$ is significantly more involved [HR08, Gei22].
To facilitate reasoning about the $k$-fold product, it would be ideal if each $\mathcal{D}_{b}$ had a computationally indistinguishable proxy distribution $\widetilde{\mathcal{D}}_{b}$ such that $d_{s}(\mathcal{D}_{0}^{\otimes k},\mathcal{D}_{1}^{\otimes k})\approx d_{\mathrm{TV}}(\widetilde{\mathcal{D}}_{0}^{\otimes k},\widetilde{\mathcal{D}}_{1}^{\otimes k})$. A recent result [MPV25] based on multicalibration [HKRR18] achieves almost exactly this, but requires a second size bound $s^{\prime}=O(sk/\varepsilon^{6})+(k/\varepsilon)^{O(1)}$ to make the connection bidirectional. Our first result in this setting improves the leading exponent on $1/\varepsilon$ from $6$ to $2$ using calibrated multiaccuracy, a weaker but more efficiently achievable condition than multicalibration that has been a theme of several recent works, as we discuss further in Section 1.4.
Theorem 1.17.
Given $s,k\in\mathbb{N}$, $\varepsilon>0$ and $\mathcal{D}_{0},\mathcal{D}_{1}$, there exist $\widetilde{\mathcal{D}}_{0},\widetilde{\mathcal{D}}_{1}$ such that $d_{s}(\mathcal{D}_{b},\widetilde{\mathcal{D}}_{b})\leq\varepsilon$ and
$d_{s}(\mathcal{D}_{0}^{\otimes k},\mathcal{D}_{1}^{\otimes k})-k\varepsilon\leq d_{\mathrm{TV}}(\widetilde{\mathcal{D}}_{0}^{\otimes k},\widetilde{\mathcal{D}}_{1}^{\otimes k})\leq d_{s^{\prime}}(\mathcal{D}_{0}^{\otimes k},\mathcal{D}_{1}^{\otimes k})+k\varepsilon,$
where $s^{\prime}=O(sk/\varepsilon^{2})+(k/\varepsilon)^{O(1)}$. One can enforce $\widetilde{\mathcal{D}}_{0}=\mathcal{D}_{0}$ with $s^{\prime}=O(sk/\varepsilon^{4})+(k/\varepsilon)^{O(1)}$.
Much like the main result of [MPV25], Theorem 1.17 is powerful because it allows one to translate statistical or information-theoretic arguments about $k$ independent draws from $\widetilde{\mathcal{D}}_{0}$ or $\widetilde{\mathcal{D}}_{1}$ to computational statements about $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$. The advantage of our theorem compared to that of [MPV25] lies in the smaller complexity gap ($1/\varepsilon^{2}$ versus $1/\varepsilon^{6}$, or $1/\varepsilon^{4}$ versus $1/\varepsilon^{12}$ for the single-proxy version), which measures the quantitative relationship between $s$ and $s^{\prime}$.
Our second result in this setting eliminates the complexity gap entirely using supersimulators, the same technical tool underlying the proof of our first result, Theorem 1.2. For the following result, we now allow ourselves to alter the distinguisher size bound $s$ within a certain interval, with the specific choice depending on the pair $\mathcal{D}_{0},\mathcal{D}_{1}$. The reason for this will become clear shortly when we discuss supersimulators in more detail.
Theorem 1.18.
Given $s,k\in\mathbb{N}$, $\varepsilon>0$, $\mathcal{D}_{0},\mathcal{D}_{1}$, there exist $s^{\prime},\widetilde{\mathcal{D}}_{0},\widetilde{\mathcal{D}}_{1}$ such that $d_{s^{\prime}}(\mathcal{D}_{b},\widetilde{\mathcal{D}}_{b})\leq\varepsilon$ and
$d_{s^{\prime}}(\mathcal{D}_{0}^{\otimes k},\mathcal{D}_{1}^{\otimes k})-k\varepsilon\leq d_{\mathrm{TV}}(\widetilde{\mathcal{D}}_{0}^{\otimes k},\widetilde{\mathcal{D}}_{1}^{\otimes k})\leq d_{s^{\prime}}(\mathcal{D}_{0}^{\otimes k},\mathcal{D}_{1}^{\otimes k})+k\varepsilon,$
where $s^{\prime}\in[s,k^{O(1/\varepsilon^{2})}s]$. One can enforce $\widetilde{\mathcal{D}}_{0}=\mathcal{D}_{0}$ with $s^{\prime}\in[s,k^{O(1/\varepsilon^{4})}s]$.
The advantage of Theorem 1.18 is that the same term $d_{s^{\prime}}(\mathcal{D}_{0}^{\otimes k},\mathcal{D}_{1}^{\otimes k})$ appears on both the left and right side of the chain of inequalities. This yields a tighter characterization in terms of $d_{\mathrm{TV}}(\widetilde{\mathcal{D}}_{0}^{\otimes k},\widetilde{\mathcal{D}}_{1}^{\otimes k})$ than in Theorem 1.17, in which distinct terms $d_{s}(\mathcal{D}_{0}^{\otimes k},\mathcal{D}_{1}^{\otimes k})\leq d_{s^{\prime}}(\mathcal{D}_{0}^{\otimes k},\mathcal{D}_{1}^{\otimes k})$, which potentially differ significantly, are used for the left and right side of the chain of inequalities. We prove both Theorems 1.17 and 1.18 in Section 5.